### v1

# Equilibrium and nonequilibrium behaviours of the Ising ferromagnetic thick cubic shell

Abstract: We have studied the equilibrium and nonequilibrium behaviours of the Ising ferromagnetic thick cubic shell by Monte Carlo simulation. Our goal is to find the dependence of the responses on the thickness of the shell. In the equilibrium results, we found that the pseudo-critical temperature of ferro-para phase transition of thick cubic shell increases with the increase of the thickness following a hyperbolic tangent relation. In the nonequilibrium studies, the relaxation time has been found to decrease with the increase of the thickness of the cubic shell. Here three different regimes are found, namely rapid fall, plateau and linear region. The metastable behaviour has been studied also as another kind of non-equilibrium response. The metastable lifetime has been studied as function of the thickness of the cubic shell. A non-monotonic variation of metastable lifetime with the thickness of the shell is observed. A specified thickness for longest-lived metastability has been identified.
Keywords: Ising ferromagnet; Monte Carlo simulation; Metropolis algorithm; Critical point; Relaxation; Metastable behaviour
PACS Nos: 05.10.Ln; 05.70.Fh; 75.10.Hk; 75.60.Jk; 64.60.My
Corresponding author
1 Introduction:
The Ising model, after a glorious 100 years of its discovery, is still a model of central attraction in the research of magnetism. The equilibrium phase transition[1, 2, 3] in Ising ferromagnet has historical importance in the field of phase transition in equilibrium statistical physics. In addition to studies on critical behaviours in regular geometric lattices, the ferro-para phase transition on the fractal lattices has revealed[4, 5, 6] many interesting results. The structure and geometry play a significant role in the critical behaviour of Ising ferromagnets.
The nonequilibrium behaviours of Ising ferromagnet have also drawn the remarkable attention of the researchers. One such nonequilibrium behaviour of Ising ferromagnet is hidden in the relaxation behaviour. In the last few decades, a notably significant number of studies have been conducted. The investigations on the decay of metastable states, domain growth and phase ordering kinetics engulfed the significant part of the literature on nonequilibrium statistical mechanics. Historically, the first relaxation dynamics was studied by Suzuki and Kubo[7] in a Glauber kinetic Ising ferromagnet using mean-field approximation. The linearized mean-field dynamical equation (near the critical temperature) can give rise to the exponential relaxation of magnetisation and also shows the critical slowing down. The relaxation behaviour of Ising ferromagnet near and away from the critical temperature has also been thoroughly investigated by extensive Monte Carlo (MC) simulation[8]. The evidence of non-exponential relaxation can be found experimentally in magnetic nanoparticle [9] for hyperthermia and manganese single-molecule magnet[10]. The Monte Carlo results of stretched exponential relaxation below the critical temperature of the large scale Ising ferromagnet has also been reported[11]. Extremely slow relaxation has been observed[12] near the first order transition of the three-dimensional Ising ferromagnet. Recently, the effects of boundary conditions, geometry, and dynamical rules have been studied[13] in a two-dimensional Ising ferromagnet by MC simulation.
Ising ferromagnetic system has been widely used over the past few decades to study nucleation [14, 15, 16, 17], domain growth kinetics [18] and metastable behaviour. The magnetic field dependences of metastable lifetime are studied [19, 20] in the Ising ferromagnet by MC simulation. The heat-assisted magnetisation reversal in ultra-thin ferromagnetic film has been studied [21] using kinetic Monte Carlo simulation, leading to enhanced nucleation. The magnetisation reversal in general spin-s Ising ferromagnetic film has been studied[22, 23] recently by employing MC simulation to find the role of anisotropy. An interesting effective two-dimensional behaviour of metastability has recently been reported[24] for a thin Ising ferromagnetic cubic shell whose Euclidean dimension is three indeed. These studies hold significance beyond just pedagogical interest. Data storage and magnetic recording [25, 26] in real-life applications require an understanding of the dynamical aspects of ferromagnetic samples, which can be visualized through MC simulations.
In the above mentioned survey, we have not found any such studies of equilibrium and nonequilibrium behaviours of the thick Ising ferromagnetic shell where the thickness may influence on the behaviours of the ferromagnet. In this manuscript we report the results of our MC investigations of the equilibrium and nonequilibrium responses of a thick Ising ferromagnetic cubic shell and studied the physical quantities as function of the thickness of the shell. The paper is organised as follows: the next section introduces the model and the simulation methodology. Section-3 contains the simulational results with data and figures. The papers ends with a summary in Section-4.
2 Model and Simulation method:
The classical Ising model is described on a simple cubic shell. The total energy of the system with uniform nearest-neighbour interaction, placed in a uniform external magnetic field $h_{\rm ext}$ is determined by the following Hamiltonian,
$H=-J\sum_{<i,j>}S_{i}S_{j}-h_{ext}\sum_{i}S_{i}$ (1)
where spin at $i^{th}$ site $S_{i}$ can take two discrete values +1 and -1 only i.e., $S_{i}\in\{-1,+1\}$; $\forall~{}i$. $J(>0)$ denotes the uniform ferromagnetic coupling between nearest-neighbour spins.
We have considered the Cubic shell of size $L$, having thickness $\Delta$ with open boundary conditions in all three directions. The system contains the total sites of $N_{s}=L^{3}-(L-2\Delta)^{3}$. A schematic cross-sectional diagram of the cubic shell is provided in Figure 1. Within the thick outer shell there exists a hollow non-magnetic cavity.
The spins in the kinetic Ising model interact with a large heat bath at a constant temperature that induces spontaneous random spin-flips (only one spin at once). Monte Carlo simulation with Metropolis single spin algorithm is employed here to study the dynamics of kinetic Ising model. Starting from our initial configuration, we have chosen a site in random manner (say $i$). At any fixed temperature ($T$) and external field ($h_{ext}$) the spin-flip ($S_{i}$) at randomly chosen site will be accepted based on transition probability determined by Metropolis rule[27].
$P(S_{i}\rightarrow-S_{i})=Min\Biggl{[}1,\exp\biggl{(}-\frac{\Delta E_{i}}{k_{B}T}\biggr{)}\Biggr{]}$ (2a)
where $\Delta E_{i}$ is the change in energy (in unit of $J$) caused by the proposed change in spin configuration at $i^{th}$ site.
$\Delta E_{i}=2S_{i}\Biggl{[}J\sum_{NN}S_{j}+h_{ext}\Biggr{]}$ (2b)
$k_{B}$ is the Boltzmann constant and $T$ is the temperature of the system measured in the unit of $J/k_{B}$. Here, $J=1$ and $k_{B}=1$ are chosen conventionally for simplicity. According to the protocol, if probability $P(S_{i}\rightarrow-S_{i})=1$ then the spin must flip. If $P(S_{i}\rightarrow-S_{i})<1$, then the spin will flip with probability $P$. An uniformly distributed (in the range $[0,1]$) random number ‘$r$’ is generated. Now, the proposed spin-flip is accepted if the condition $r\leq P(S_{i}\rightarrow-S_{i}$ is satisfied. Otherwise, the proposal of spin-flip will be rejected and the spin will retain its original state. This randomly selected $N_{s}$ number of updates constitutes one Monte Carlo Step per Site (MCSS) denoted as unit of time ($t$) used in our simulation.
Instantaneous magnetisation is recorded at each MCSS,
$m(t)=\frac{1}{N_{s}}\sum_{i}S{{}_{i}}.$ (3)
3 Results:
In our present study we have thoroughly analysed the continuous (ferro to para) phase transition in the Ising cubic shell with $L=50$, focusing on how the critical point shifts with changes in shell thickness $\Delta$. An approximate estimation of $T_{c}$ is necessary to maintain the equivalent thermal condition across the systems with different shell thickness, ensuring that the thermal conditions remain comparable throughout the analysis. Transient behaviours for example, relaxation and reversal of magnetisation, are also investigated for Ising cubic shell with varying thickness $\Delta$.
3.1 Ferromagnetic Phase transition:
In order to study the continuous ferro-para phase transition, the Ising ferromagnetic thick cubic shell with $L=50$ and thickness $\Delta$ is gradually cooled down in small temperature steps of $\delta T=0.05$, starting from high temperature random spin configuration (paramagnetic phase). At each time (MCSS) step, the corresponding thermodynamic quantities of the system are measured. The instantaneous magnetisation $m(t)$ has been calculated. The equilibrium magnetisation in the absence of external field ($h_{\rm ext}=0$) is defined as $M=\langle m\rangle$, while susceptibility is obtained using $\chi=\frac{N_{s}}{k_{B}T}(\langle m^{2}\rangle-{\langle m\rangle}^{2})$. Here .$\langle$.$\rangle$ . . ⟩ denotes the time average of thermodynamic quantities, approximately identical to the ensemble average in the ergodic limit. Thermodynamic quantities are recorded and averaged over $10^{4}$ MCSS to ensure more accurate and stable measurements. As an example, for a cubic shell with $L=50$ and $\Delta=5$, the total length of the simulation is 75000 MCSS. Out of which initial 65000 MCSS are discarded to ensure the ergodic limit and the statistical measurements are averaged over the rest $10^{4}$ MCSS. The time-averaged thermodynamic quantities are further averaged over 25-50 random samples depending on the system’s size and thickness to enhance statistical accuracy.
We have thoroughly observed the variation of magnetisation ($M$) and susceptibility ($\chi$) with temperature ($T$) and graphically demonstrated in the Figure 2 for different values of thickness $\Delta$ of the Ising cubic shell system. At high temperature, the order parameter i.e., magnetisation ($M$) of the system is almost zero, ensuring the paramagnetic phase. As we decrease the temperature, the magnetisation becomes non-zero at the transition temperature (or pseudo-critical temperature $T_{c}^{p}$). This is depicted in Figure 2(a). The susceptibility of the system exhibits a sharp peak at the transition temperature ($T_{c}^{p}$). The position of the peak of the susceptibility, as shown in Figure 2(b) provides an estimate of pseudo-critical temperature ($T_{c}^{p}$) of the Ising cubic shell for various thickness values $\Delta=$1, 2, 3, 5, 10, 15, 20, 22 and 25.
The transition temperature of an Ising ferromagnetic cubic shell with $L=50$ and thickness $\Delta=1$ is estimated to be $T_{c}^{p}=2.30\pm 0.05$, remarkably close to Onsager’s exact value of $T_{c}=2.269$ (in $J/k_{B}$ unit) for a two-dimensional Ising system, which suggests that the ferro-para phase transition in the thinnest cubic shell (monolayer) closely resembles that of a flat two-dimensional system. Our results show that the transition temperature increases with the increase in the thickness of the cubic shell. This upward shift in $T_{c}^{p}$ continues up to a saturating value of $\Delta$ (around $\Delta=15$ here), indicating a direct dependence of pseudo-critical temperature on the shell-thickness before it stabilizes at higher values of $\Delta$. These observations lead us to the conclusion that beyond the saturating value of shell-thickness ($\Delta$ larger than 15) the presence of non-magnetic cavity within the thick shell of Ising spins does not influence the equilibrium ferro-para phase transition. Figure 3 presents the pseudo-critical temperature ($T_{c}^{p}$) as a function of the cubic shell’s thickness $\Delta$. As seen from Figure 3, the pseudo-critical temperature $T_{c}^{p}$ increases with thickness and after reaching a saturating value of $\Delta$, $T_{c}^{p}$ approaches Monte Carlo estimated value for 3D Ising ferromagnetic system $T_{c}=4.511~{}J/k_{B}$. The variation of $T_{c}^{p}$ with thickness $\Delta$ is well described by the function $T_{c}^{p}(\Delta)=a\times\tanh(b~{}\Delta)+c$, with best-fit parameters $a=2.98\pm 0.19$, $b=0.321\pm 0.026$ and $c=1.43\pm 0.19$. This fit, represented by the solid green line in Figure 3 yields the value of statistical $\chi^{2}=0.046$ with degrees of freedom = 6. This finding is significant as it provides new insights into the equilibrium behaviour of Ising ferromagnetic thick cubic shells and has not been reported elsewhere.
3.2 Finite size analysis:
In our previous study[24], we have roughly estimated the pseudo-critical temperature $T_{c}^{p}$ of the Ising ferromagnetic cubic shell with thickness $\Delta=1$ for a finite-sized system of $L=100$. According to our findings, the thermal variation of susceptibility produces a sharp peak at $T_{c}^{p}=2.28J/k_{B}$. We know that the pseudo-critical temperature converges to the true critical temperature in the thermodynamic limit $L\rightarrow\infty$, where both susceptibility and correlation length diverge at the critical point of the ferro-para phase transition.
We have carried out a systematic finite size analysis of Ising ferromagnetic cubic shell with $\Delta=1$ for three different values of $L=75,~{}100~{}\rm and~{}150$ and investigated the thermal variation of magnetisation($M$), susceptibility($\chi$) and fourth order Binder cumulant $U_{L}=1-\frac{\langle m^{4}\rangle}{3{\langle m^{2}\rangle}^{2}}$, as depicted in Figure 4 (a), (b) and (c), respectively. From Figure 4(b) it is evident that the height of the maximum susceptibility gradually increases with $L$, symbolising the growth of the critical correlation. This indicates that the susceptibility of the system will eventually diverge at critical temperature for $L\rightarrow\infty$. Figure 4 (c) illustrates that the Binder cumulant $U_{L}$ for different values of $L$ intersects in $T_{c}$. Unfortunately, we could not provide a clear intersection point of $U_{L}$ curves, approximately around $T=2.26$. Although our results are incapable of having any scaling behaviour (and critical exponents), the divergence of susceptibility in the thermodynamic limit is evident.
3.3 Magnetic Relaxation behaviour:
The ferromagnetic system is initially perturbed from its equilibrium state by applying a very strong magnetic field, forcing it into a perfectly ordered configuration with magnetisation $m=1$. When the external perturbation (strong magnetic field) is abruptly removed, the system relaxes back to its steady state of equilibrium through the relaxation of magnetisation.
In this study, we have investigated the magnetic relaxation in the Ising ferromagnetic cubic shell. We have started from an initial state where all spins are aligned upwards (as if the system is in a strong magnetic field). This corresponds to a state of maximum magnetisation and certainly not the equilibrium state at any temperature in paramagnetic phase. Now, immediately after the removal of the strong field, the magnetisation of the system decays in time until it reaches equilibrium (with $m(t)$ vanishingly small) at any finite temperature $T$ (above $T_{c}^{p}$). The temperature of the system is maintained in the paramagnetic phase, specifically at $T=1.10~{}T_{c}^{p}(\Delta)$. It may be noted here that this ($T=1.10~{}T_{c}^{p}(\Delta)$) condition serves as an equivalent thermal condition for all values of shell thickness ($\Delta$).
Here, Figure 5 illustrates the time evolution or the relaxation of magnetisation for different thickness $\Delta$ of the Ising ferromagnetic cubic shell of $L=50$, presented in the linear scale (5(a))and semi-log scale (5(b)), respectively. The results are obtained, averaging over 8000-25000 random samples, depending on the system size to ensure statistical accuracy. The semi-logarithmic plot of decay of the magnetisation, as shown in Figure 6, clearly exhibits the exponential nature of magnetic relaxation, of the following form,
$m(t)\sim\exp(-t/\tau_{\rm relax})$ (4)
Where $\tau_{\rm relax}$ is the characteristic time associated with magnetic relaxation called relaxation time. The data of magnetisation decay $m(t)$ have been analysed by fitting to an exponential function $f(t)=a\exp(-b~{}t/10^{3})$, where $a$ and $b$ are fitting parameters corresponding to the initial magnetization amplitude and decay rate, respectively. The solid lines in Figure 6 represent the exponential fit for $m(t)$ on semi-logarithmic scale. The inverse relaxation time $1/\tau_{\rm relax}$ is derived from the fitted parameter $b$ of the function $f(t)$. The best-fit values of the fitting parameters are provided in Table 1.
Table 1: Best fitted parameters (Cubic shell) corresponding to Figure 6.
Thickness $\Delta$ Slope (b) ×$\times 10^{-3}$ $\chi^{2}$ DOF
1 $10.032\pm 0.008$ 0.00082 489
2 $17.061\pm 0.012$ 0.00022 385
5 $27.823\pm 0.042$ 0.00023 229
20 $34.571\pm 0.139$ $4.83\times 10^{-5}$ 201
25 $50.319\pm 0.083$ $1.94\times 10^{-5}$ 114
We have systematically recorded the relaxation time $\tau_{\rm relax}$ of the Ising ferromagnetic cubic shell by varying its shell-thickness $\Delta$ at two different temperatures, say $T=1.10T_{c}^{p}(\Delta)$ and $1.15T_{c}^{p}(\Delta)$. The dependence of $\tau_{\rm relax}$ on thickness $\Delta$ is depicted in Figure 7. The obtained results reveal that the relaxation time decreases as we increase the shell-thickness. In the beginning, the relaxation time $\tau_{\rm relax}$ exhibits a rapid decrease with shell-thickness $\Delta$, suggesting that thinner shells experience more prolonged relaxation processes. This rapid fall region is extended within a range of values of thickness $\Delta$ (say $\Delta=1-6$). As $\Delta$ continues to increase, the relaxation time approaches a near-saturation point for intermediate values of $\Delta$ (say $\Delta=10-20$), indicating a regime where the influence of shell-thickness on relaxation is insignificant. In that plateau region, the relaxation dynamics become less sensitive to the existence of internal non-magnetic cavity. Beyond this intermediate regime ($\Delta>20$), $\tau_{\rm relax}$ resumes a linear decrease with increasing $\Delta$, highlighting a renewed sensitivity of relaxation dynamics on the shell thickness at larger values. Our findings indicate a complex interplay between magnetic relaxation and shell-thickness. This is interestingly new result.
3.4 Metastable behaviour:
The ferromagnetic system, below the critical temperature, shows metastable behaviour. If the ferromagnetic system is prepared initially with all spins up and allowed to relax in the presence of a small magnetic field (in the opposite direction), it retains its metastable phase during a period. This is called metastable lifetime. After this, the system jumps to its stable state of negative magnetisation. The minimum time required by the system to gain negative magnetisation is defined as reversal time or metastable lifetime $\tau_{\rm rev}$.
In this section, we provide a comprehensive analysis of the magnetisation reversal in the Ising ferromagnetic cubic shell, focusing on the effects of shell-thickness on the metastable behaviour. The Ising ferromagnetic shell of $L=50$, initially configured with all spins up i.e., $S_{i}=+1,~{}\forall i$ condition, has been exposed to a weak negative magnetic field $h_{\rm ext}=-0.20$. Temperature of the system is fixed at $T=0.8T_{c}^{p}$ (ferromagnetic phase). This choice is arbitrary and represents the equivalent thermal condition for the cubic shell of various shell-thickness. We have critically examined the the decay of metastable state of a single sample and illustrated in Figure 8, showing the system’s metastable behaviour for various shell thickness $\Delta=$ 1, 3, 15 and 25. These findings offer valuable insight into how the thickness of the shell influences the lifetime of the metastable state.
In our analysis, we have referred to the metastable lifetime $\tau_{\rm meta}$ as the minimum time steps required for the magnetisation to drop below $m=0.7$. This method of defining the metastable lifetime is used in previous studies [20]. Similarly, the reversal time[22] denoted as $\tau_{\rm rev}$ refers to the time by which the magnetisation crosses $m=0$ and changes sign. To obtain statistically significant data, we have calculated the mean value of $\tau_{\rm rev}$ as well as $\tau_{\rm meta}$ by averaging over 1000-5000 random ferromagnetic samples.
The variation of mean $\tau_{\rm meta}$ with thickness $\Delta$ is graphically represented in Figure 9 (a) for three distinct external field strengths $h_{\rm ext}=-0.15,~{}-0.20~{}\rm and~{}-0.25$. Our findings reveal that the mean metastable lifetime $\tau_{\rm meta}$ changes non-monotonically with increasing shell thickness $\Delta$. Furthermore, we have studied the mean reversal time $\tau_{\rm rev}$ as a function of thickness $\Delta$ for the same values of field strength, as shown in Figure 9 (b). Both of these two time scales $\tau_{\rm meta}$ and $\tau_{\rm rev}$ exhibit almost a similar kind of non-monotonic variation with $\Delta$. The mean reversal time $\tau_{\rm rev}$ is observed to be at its minimum for the thinnest shell ($\Delta=1$), indicating a relatively faster reversal process. Interestingly, a specific range of shell-thickness ($\Delta=3-5$) is observed which maximizes the time scale ($\tau_{\rm meta}$ and $\tau_{\rm rev}$). A particular thickness within this range leads to significantly prolonged metastability. Notably, no similar results have been reported elsewhere.
4 Summary:
In this manuscript, we have reported the results of both equilibrium and nonequilibrium behaviour of a thick cubic Ising ferromagnetic shell. The results are obtained using Monte Carlo simulation. In the equilibrium case, the ferro-para phase transition is studied. The transition temperature (pseudo-critical temperature $T_{c}^{p}$) is obtained from the temperature which maximizes the susceptibility. The pseudo-critical temperature has been found to be dependent on the thickness ($\Delta$) of the cubic shell. Our numerical results show that the pseudo-critical temperature increases as the thickness of the cubic shell increases. For the thinnest shell (monolayer) the pseudo-critical temperature is close to the transition temperature of a two-dimensional Ising ferromagnet. On the other hand, for maximum thickness (where the thick cubic shell becomes a solid cube), the pseudo-critical temperature is approximately equal to the Monte Carlo estimate of the ferro-para transition temperature of a solid cube. In the intermediate region, the variation of the pseudo-critical temperature with the thickness of the shell is shown with a best fitted function like $T_{c}^{p}(\Delta)=a\times\tanh(b~{}\Delta)+c$, where the values of $a$, $b$ and $c$ are $a=2.98\pm 0.19$, $b=0.321\pm 0.026$ and $c=1.43\pm 0.19$ estimated for the best fit. This is an interesting result and has not been found before in the literature. We have also studied the finite size analysis. However, our results are too imprecise to identify any scaling behaviour with critical exponents.
In the nonequilibrium case, we have studied the (i) relaxation behaviour and the (ii) metastable behaviour of the thick cubic shell. In the relaxation behaviour, we have allowed the system to relax from a state where all spins are up. The state of all spin-up is the equilibrium state (ground state) for $T=0$. However, an equilibrium state at any finite temperature would be achieved through the relaxation of magnetisation. The equivalent thermal condition is prepared as $T=f\times T_{c}^{P}(\Delta)$. We have used $f=1.10~{}\rm and~{}1.15$ for all values of the thickness ($\Delta$) of the shell. The exponential relaxation is observed for all values of $\Delta$. However, as the shell-thickness ($\Delta$) increases, the relaxation time $\tau_{relax}$ is found to decrease. It may be noted here that for smaller values of thickness the relaxation time decreases rapidly, in an intermediate range of the values of the thickness we have observed a plateau (where the relaxation time is almost independent of thickness) region. For higher values of thickness the relaxation time linearly decreases with the thickness. As far as we are aware, we have not encountered similar studies or results in the literature.
The decay of metastable state is another example of non-equilibrium behaviour. The Ising ferromagnet shows metastability and the metastable lifetime depends on the magnitude of the applied magnetic field and it is well studied. In this paper, we have studied the dependences of metastable lifetime on the thickness of the cubic shell of Ising ferromagnet. Interestingly, the metastable lifetime also shows a non-monotonic variation with the thickness of the cubic shell. Initially, as the value of shell-thickness increases, the metastable lifetime increases and after a specific value of the shell thickness the metastable lifetime decreases. There exists a specific value of the thickness, for which the metastable lifetime becomes maximum. These results are not found in previous reports.
Our study may trigger the interest of the experimentalist to study these kinds of behaviours in real ferromagnetic sample having cavity. Depending on the size of the cavity (non-magnetic), the transition temperature and the metastable lifetime can be regulated by varying the size of the cavity inside the ferromagnetic sample.
Acknowledgements: IT acknowledges UGC JRF, Govt. of India for financial support. We are thankful to the computational facilities provided by Presidency University, Kolkata. Data availability statement: Data will be available on reasonable request to Ishita Tikader.
Conflict of interest statement: We declare that this manuscript is free from any conflict of interest.
Funding statement: No funding was received particularly to support this work.
Authors’ contributions: Ishita Tikader- Developed the code, prepared the figures, and wrote the manuscript. Muktish Acharyya-Conceptualized the problem, analysed the results and wrote the manuscript.